OpenStudy (anonymous):

42.6 g Cu are combined with 84.0 g of HNO3 according to the reaction: 3 Cu + 8 HNO3 → 3 Cu(NO3)2 + 2 NO + 4 H2O. Which reagent is limiting and how many grams of Cu(NO3)2 are produced?

OpenStudy (anonymous):

HNO3 limiting reagent and 93.76 g are produced

OpenStudy (anonymous):

Thanks how did you get that?

OpenStudy (anonymous):

wait one sec

OpenStudy (anonymous):

42.6gCU x 1 mole CU / 63.55gCU = .67 mole CU available 84.0gHNO3 x 1mole HNO3 / 63.02gHNO3 = 1.33 mole HNO3 available .67 mol CU x 8 mole HNO3(needed) / 3 mole CU =1.78 mole HNO3 needed and only 1.33 mole available therefore HNO3 is Limiting. 84g HNO3 x 1 mole HNO3 / 63.02 HNO3 x 3 moles Cu(NO3)2 / 8 mole HNO3 x 187.57 g Cu(NO3)2 / 1 mole CU(NO3)2 =93.755 g of Cu(NO3)2 produced spaced it al out so it's easier to read

OpenStudy (anonymous):

Wow this process is very long...gotta study more! Thanks!

OpenStudy (chmvijay):

@nwynter now study more OK!!!!

OpenStudy (anonymous):

bleh is the equivalent of math and chem combined good luck